OpenStudy (anonymous):

Type the coordinates of the images P,Q,R,S, and T for each of the following points under the translation defined by (x,y)>(x+9,y-6). P(0,0)>P= Q(-1,1)>Q= R(-9,-3)>R= S(7,14)>S= T(r,s)>T=

OpenStudy (anonymous):

can you please help me. I'm lost on how to do this

OpenStudy (anonymous):

P=9,-6 Q=8,-5 R=0,-9 S=16,8 T=r+9, s-6